What is an electrical resonance ? State it's real daily life application
1
Expert's answer
2021-07-26T11:26:01-0400

1. electrical resonance is a state when inductive effect of a circuit is nullified by the capacitive effect and the circuit, then, behaves as a pure resistance.


2.This happens because inductance and capacitance have opposite properties - while inductive reactance increases with frequency, capacitive reactance decreases with frequency.


LC circuit

impedence

"z=\sqrt{R^2+(X_L-X_c)^2}"

Z=0

Or Z=infinite

In a simple series LC circuit

resonance occurs

when

inductive reactance= capacitive reactance

"X_L=X_c"

giving zero impedance and infinite current.


Example LC circuit

Resources frequency

"w_r=\frac{1}{\sqrt{LC}}"

"f_r=\frac{1}{2\pi\sqrt{LC}}"

1. Pendulum

2. Microwave oven

3. Radio

4. Swing

5. Guitar